3. You titrate 125 mL of a solution of ammonia (NH3) with 1.0 M HClO4. The equivalence point is reached after adding 25 mL of the titrant.

a.) What kind of titration was this (WA/SB; WB/SA; SA/SB; polyprotic A (or B) /SB (or SA); WA/WB)?

b.) Write a balanced reaction for the titration.

c.) What is the concentration of the NH3?

d.) Would the resulting solution be acidic, basic, or neutral?

4. Write the Keq expression for the following equilibria:

a.) H3PO4 (aq) + H2O (l) ⇌ H3O + (aq) + H2PO4 - (aq)

b.) HCO3 - (aq) + H+ (aq) ⇌ H2O (l) + H2CO3 (aq)

c.) PbSO4 (s) ⇌ Pb+2 (aq) + SO4 -2 (aq)

6. You’ve been tasked with disposing of some radioactive waste from the manufacture of medical imaging agents. The main radioactive isotope in this waste is 131I (which has a half-life of ~8.0 days and decays primarily via a β - mechanism). There are companies that specialize in the disposal of radioisotopes, but they are very expensive. The local landfill has agreed to allow you to dispose of the waste as long as the radioactivity is below 1 x 10-3 mCi.

a.) If the waste originally has an activity of 10 Ci, how long would you have to store the waste to be able to dispose of it in the landfill? (Remember to show your work for full credit)

b.) What kind of container should you store the waste in while it’s decaying?
Answer #1

3)

a) strong acid weak base so WB/SA

b) NH3 + HClO4 ----> NH4+ + ClO4- ammonium perchlorate is the product

c)conc of ammonia is 0.2 M 25*1=Mbase*125

d)solution is acidic(ammonium perchlorate)

4) a)keq=[H3O +][H2PO4 - ]/[H3PO4] b) keq=[H2CO3]/[H+] [HCO3 -] c)keq=[Pb+2] [SO4 -2]

5) a) assuming first order kinetics

k=ln2/half life period half life=8 days=691200s so k=1*10^(-6)

A=A0EXP(-kt) A=activity at any time=1*10^(-6) ci (threshold for disposal) ,A0=initial activity=10 ci

10^(-7)=EXP(-kt)

taking log on both sides    

-7*2.303=-(1*10^(-6))*t

so t=16121000s=186.58 days

b) use a thick aluminium container which can protect beta radiation from coming out